A woman driving a car 14 ft long is passing a truck 30 ft long. The truck is

traveling at 50 mi/h.

How fast must the woman drive her car so that she can pass the truck

completely in 6 seconds, 


1
Expert's answer
2021-09-06T11:35:12-0400

"30\ ft=9.144\ m"

"14\ ft=4.2672\ m"


"50\ mi/h=22.352\ m/s"


"(v_2-v_1)t=s_1+s_2\to(v_2-22.352)\cdot6=9.144+4.2672\to"


"v_2=24.5872\ (m/s)=55\ (mi/h)" . Answer